OpenStudy (anonymous):

Simplify the expression: (4xy)-2

OpenStudy (unklerhaukus):

\[(4xy)-2=2(2xy-1)\]

OpenStudy (anonymous):

would the answer be 16x squared y squared

OpenStudy (unklerhaukus):

\[(4xy)^{-2}=\left(\frac{1}{16x^2y^2}\right)^{-1}=16x^2y^2\]